OpenStudy (anonymous):

An equilateral triangle has a perimeter of 21 m. If the area is multiplied by 1/2 (one-half), what is the new side length?

OpenStudy (anonymous):

A = 1/4(s^2)(√3) A = 1/4(49)(√3) A = 49√3/4 After multiplying by 1/2, then it's 49√3/8 1/4(s^2)(√3) = 49√3/8 s^2 = 49/2 s = 7/√2 = 7√2/2